Given the equations 3x + y = 17, 5y + z = 14 and 3x + 5z = 41, what is the value of the sum x + y + z?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

Given the equations 3x + y = 17, 5y + z = 14 and 3x + 5z = 41, what is the value of the sum x + y + z?
3x + y = 17, ---------------1
5y + z = 14 ----------------2
3x + 5z = 41, ---------------3
Multiply eq1 by5 and subtract 2 from it
15x+5y=85
5y+z=14
15x-z=71
Multiply eq 3 by 5 and solve with the above
15x+25z=205
-26z=- 134
Z= 5.15
Y=1.77
X= 5.08

Answer by richwmiller(17219)

total = 12, x = 66/13, y = 23/13, z = 67/13